# AIDE: Antithetical, Intent-based, and Diverse Example-Based Explanations

### v1

@@ Introduction
Failure of ML-based systems in numerous cases, e.g., due to data errors, biases, misalignment (Osoba and Welser IV 2017; Tashea 2017; Seymour et al. 2023; Burema et al. 2023), has prompted researchers to work on explainability techniques. Different taxonomies for such methods exist, e.g., (Guidotti et al. 2018), but one common classification is on the type of explanation generated (Molnar 2022). Model-based methods involve creating interpretable surrogate models, such as decision trees or linear models, which approximate the complex black box ML model (Ribeiro, Singh, and Guestrin 2018; Silva et al. 2020). Feature-based methods focus on pinpointing important features of the input, such as words in text or parts in an image, which contribute the most to the prediction (Ribeiro, Singh, and Guestrin 2016; Fong, Patrick, and Vedaldi 2019; Ancona 2017). Example-based methods provide explanations for a specific target outcome by deriving the importance of training samples (Ilyas et al. 2022; Garima et al. 2020; Koh and Liang 2017; Kwon and Zou 2022; Ghorbani and Zou 2019; Park et al. 2023), or provide a global overview of the model identifying representative examples (Yeh et al. 2018; Pruthi et al. 2020).
Example-based explainability offers several advantages. They are typically model-agnostic, and offer easy to understand explanations. More importantly, as they seek to discover a causal relationship between training examples and model behavior, they can assist in model debugging and data cleansing (Hara, Nitanda, and Maehara 2019). However, they have two key limitations.
First, they don’t offer contrastivity (Nauta et al. 2023), which is key aspect in how humans understand decisions (Lipton 1990). While most methods can distinguish between supporters (aka proponents, helpful or excitatory examples), and opposers (aka opponents, harmful or inhibitory examples), they do not relate this information to ground truth labels (examples of class same as or different than predicted) or to the explanation intent (is the prediction correct/wrong, hard to tell). Contrastivity is the hallmark feature of counterfactual explanations (Wachter, Mittelstadt, and Russell 2017) and a major part of their appeal. Even though counterfactuals are instances, these are by design imaginary, not necessarily plausible (Pawelczyk, Broelemann, and Kasneci 2020) or robust (Slack et al. 2021). In essence, counterfactual explanations offer feature-based explainability, revealing the important feature values contributing to the outcome.
@@
In this paper, we propose a novel Antithetical, Intent-based, and Diverse Example-based (AIDE) explainability method, that offers contrastivity and is robust to outliers. At its core, AIDE is based on the concept of influence functions (Hampel 1974; Koh and Liang 2017). For a fixed target instance, the influence of a training sample is a score conveying its impact on the classifier’s outcome. Ideally, the influence is the change observed in the loss value for the target if the training sample was excluded from the training data. While influence scores can be estimated by methods, such as TraceIn (Garima et al. 2020) and Datamodels (Ilyas et al. 2022), we use the framework of the influence function approach (Koh and Liang 2017), termed IF, to efficiently compute influence scores.
To better understand AIDE’s contribution, we first showcase the issues that plague example-based explainability methods, taking IF as the representative—extensive qualitative and quantitative comparison with other methods is presented in the evaluation section. Consider a classifier that predicts whether short text messages are spam. Figure 1 shows that for the depicted target message, the prediction is spam. This is a correct prediction, and IF identifies the four most influential training samples at the top of Figure 1. We observe that explanations lack diversity, as they are highly similar to each other. More importantly, however, they lack contrastivity, as the user does not gain any insight about how the model decides what is spam and what not; all the user learns is that similar texts were labelled spam. The issue of susceptibility to outliers does not manifest in this example, mainly because the prediction is clearly correct. However, it manifests when for example what the correct prediction should be is not clear as in Figure 6.
@@
AIDE is intent-aware. A critical review (Keane et al. 2021) identifies misalignments between the design goals of XAI methods and the psychological and cognitive aspects of explainability. A key limitation they identify is that the majority of XAI methods neglect user intent during design and evaluation. Although mostly referring to counterfactual explanations, the authors highlight the importance of considering the user’s needs and goals while generating explanations. Moreover, the authors identify another common limitation in that explanations may not be plausible, which they define as not being relevant to the prediction being explained. AIDE explicitly addresses both limitations by offering intent-aware and relevant example-based explanations.
@@
AIDE outperforms state-of-the-art example-based methods. We perform an extensive quantitative and qualitative comparison against state-of-the-art methods for example-based explainability. Shapley-based approaches (Ghorbani and Zou 2019; Kwon and Zou 2022) were excluded as (a) they primarily aim to capture the overall contribution of training samples to the trained model (data valuation), and not geared for local explanations, and (b) they can be impractical for local explainability due to their high computational cost. The main conclusions drawn are as follows. Datamodels (Ilyas et al. 2022) approximate well the influence scores, but perform poorly for outlier targets. The principal reason is that Datamodels explain a class of models, and not a particular model. They thus fail to identify nuances picked up by a single model; while an unambiguous target will receive similar predictions by all models in the class, models will greatly differ in their predictions for an ambiguous target. TraceIn (Garima et al. 2020) is highly susceptible to outliers in the training data and performs poorly in tests of correctness and truthfulness. The reason lies in the way TraceIn estimates influence: it considers the difference in the total (training) loss when a training sample is included or not during checkpoints; outliers have high individual loss, contributing significantly to the total loss, and are thus awarded high importance. Regular Influence functions (IF) are similarly affected by outliers in the training data. RelatIF (Barshan, Brunet, and Dziugaite 2020) seeks to address this problem, by penalizing samples that have high loss. However, these high-loss samples can at times be highly informative. For example, to explain a target instance that is ambiguous, it is often insightful to present those outlier training examples that are similar, so as to potentially uncover interesting labelling rules or protocols. In contrast, AIDE considers outliers as long as they are relevant to the target instance.
@@
Another method similar in principle to IF is TraceIn (Garima et al. 2020) that measures the influence of a training sample $X$ on a specific test sample $X_{0}$ as the cumulative loss change on $X_{0}$ due to updates from mini-batches containing $X$. They practically approximate this with TraceInCP, which considers checkpoints during training and sums the dot product of gradients at $X$ and $X_{0}$ at each checkpoint. Another interesting and unconventional work (Ilyas et al. 2022) fixes a test point to explain and samples a large number of subsets from the training set and trains models with each of these subsets. It then trains a linear model where the input will be $1_{S}i$ encoding of a subset and the output is the performance of the model trained on this subset for the test sample of interest. The weights of the linear model will represent the importance score of a training sample in the same position. To obtain a good result a huge number of intermediate models has to be trained on subsets, which is exhaustive, and thus a faster version of datamodels was proposed by (Park et al. 2023) and claimed to preserve almost the same accuracy. However, since our focus is on the effectiveness of explanation we still use the original datamodels as a baseline.
@@
Evaluating Explanations. While significant progress has been in developing explainability methods, there is a lack of standardized metrics for evaluating their effectiveness. The authors of (Doshi-Velez and Kim 2017) distinguish three types of evaluation strategies: application-grounded, human-grounded, and functionality-grounded. A profound study of functionality-grounded strategies by (Nauta et al. 2023), advocates twelve quantifiable properties that can be evaluated to assess the quality of explanations. They categorize the state-of-the-art metrics into twelve classes depending on which property the metric focuses on and what type of explanation is provided. The following properties are most relevant for local, example-based explanations: (1) Consistency and continuity both describe how deterministic the explanation is concerning identical and similar samples, assuming that these samples should have identical and similar explanations. In many works, this aspect is also referred to as the faithfulness (Jacovi and Goldberg 2020; Adebayo et al. 2018) of explanation and has gained popularity in the explainability domain. (2) Contrastivity is the ability of an explanation to interpret classes different than the prediction class. (3) Compactness is encoded in the size of an explanation as well as calculating a redundancy in the explanation. (4) Context describes how relevant the explanation is to the user needs. (5) Controlled synthetic Data check—Controlled Experiment: a synthetic dataset is developed with predetermined reasoning, ensuring that the predictive model aligns with this reasoning, as verified through metrics like accuracy. An assessment is done to check whether the explanation provided by the model corresponds to the same reasoning embedded in the data generation process, (Adebayo et al. 2020; Chen et al. 2018). In another work (Mothilal, Sharma, and Tan 2020), two general metrics for example-based explainability are proposed, diversity, and proximity. Whereas diversity can be attached to the compactness property noting that it prevents redundancy in the explanation, proximity is a measure of how close or relevant is the explanation to the test sample. By design, AIDE offers diverse and proximal example-based explanations, which are faithful and correct as our quantitative evaluation reveals.
@@
It can be shown (Cook and Weisberg 1982) that under certain conditions (second order differentiability and convexity of the loss function) the derivative of ${\bm{\theta}}^{*}$ can be expressed as:
@@
where ${\bm{H}}_{{\bm{\theta}}^{*}}$ is the Hessian matrix (containing the second order partial derivatives) of the objective $L({\mathcal{S}},{\bm{\theta}}^{*})$ calculated at ${\bm{\theta}}^{*}={\bm{\theta}}^{*}_{0}$.
@@
To explain the prediction for ${\bm{z}}_{t}$, we use Equation 5 to compute the influence of each training example ${\bm{z}}$, which can be done efficiently as suggested in (Koh and Liang 2017). The IF explanation (Koh and Liang 2017) for the prediction for ${\bm{z}}_{t}$ consists of the top-$k$ training examples with the highest influence.
@@
Existing approaches for influence-based explainability (Koh and Liang 2017; Barshan, Brunet, and Dziugaite 2020) compile an explanation as a set of highly influential training examples. We claim that other aspects, besides high influence, are also important. Specifically, AIDE creates explanations that contain training examples with negative influence, considers their labels, their proximity to the test instance, and their diversity.
@@
Proximity. Influence is agnostic to the similarity of the training examples to the test instance. As noted (Barshan, Brunet, and Dziugaite 2020), there may exist outliers and mislabeled training examples that can exhibit high magnitude influence scores. Such examples are often globally influential, i.e., they are influential for many test instances, just because they are unusual. These are rarely useful as an explanation, and (Barshan, Brunet, and Dziugaite 2020) proposes to normalize the influence of an example with their global influence. Nonetheless, in certain cases these outliers are extremely useful, e.g., when explaining another outlier.
@@
Diversity. Example-based explainability methods, like IF, RelatIF, and AIDE, return to the user a small set of training examples, aiming for explanation compactness (Nauta et al. 2023). It is thus important that the set of examples avoids redundancy. AIDE, in contrast to prior work (Koh and Liang 2017; Barshan, Brunet, and Dziugaite 2020), considers the diversity of the explanation set. Assuming an internal representation of training examples and an appropriate similarity measure $\operatorname{sim}$, we define diversity of a set ${\mathcal{E}}$ of training examples as $D({\mathcal{E}})=1-\frac{1}{|{\mathcal{E}}|(|{\mathcal{E}}|-1)}\sum_{{\bm{z}},{\bm{z}}^{\prime}\in{\mathcal{E}}}\operatorname{sim}(\hat{{\bm{x}}},\hat{{\bm{x}}^{\prime}})$.
@@
Support. It comprises examples with positive influence and with the same label as the test instance. They play a positive role in the prediction and resemble the test instance in terms of their characteristics: “You get the same outcome with these”.
@@
To select the appropriate examples for each quadrant, we perform a series of steps. After partitioning the training examples in the four quadrants, we select only examples with high magnitude. We use the Interquartile Range (IQR) method, (Agresti and Franklin 2005), to keep examples with positive influence above $Q_{3}+\lambda IQR$, and to keep examples with negative influence below $Q_{1}-\lambda IQR$, where $Q_{1}$ and $Q_{3}$ are the first and the third quartiles of the influence distribution, $IQR=Q_{3}-Q_{1}$, and $\lambda$ is a coefficient that controls the number of high-magnitude influential points, and is empirically determined. After this filtering, we end up with a candidate set ${\mathcal{S}}_{q}$ of training examples for each quadrant $q\in\{1,2,3,4\}$.
@@
In our experiments, we used two datasets: the SMS Spam dataset [^2] https://www.kaggle.com/datasets/uciml/sms-spam-collection-dataset, which comprises a collection of text messages labeled as either spam or non-spam (ham), commonly used for text classification and a derivative dataset with pictures of dogs and fish extracted from Imagenet [^3] https://www.image-net.org/. For the spam classification task, we employed the BERT-base pre-trained word embedding model and incorporated two sequential layers to capture the specific characteristics of our data. Regarding the image classification task, we utilized a pre-trained InceptionV3 model removing the output layer and appending sequential layers to learn the peculiarity of our task. All the baselines were implemented with instructions given in their papers and GitHub repositories. The coefficient for IQR was set to $\lambda=3$ in all cases. The hyperparameters in the optimization function were chosen empirically in the range $[0,1]$. We observed that the diversity weight $\gamma$ does not affect the quality of the explanations that much, as long as it was nonzero; we set it to $\gamma=0.5$ in all experiments. The other two hyperparameters control the presence of outliers in the explanations; higher values of $\beta$ suppress outliers by giving more weight to training examples that are similar and have high enough influence. We settled to $\alpha=0.2$ and $\beta=0.8$ for all experiments. More information on reproducibility with the link to the GitHub repository can be found in the appendix. The baseline methods that we will compare AIDE to are IF (Koh and Liang 2017), RelatIF (Barshan, Brunet, and Dziugaite 2020), Datamodels (Ilyas et al. 2022), and TraceIn (Garima et al. 2020).
@@
Figure 2: Continuity in terms of explanation similarity vs. instance pair similarity.

### v2

@@ Introduction
Failure of ML-based systems in numerous cases, e.g., due to data errors, biases, misalignment (Osoba and Welser IV 2017; Tashea 2017; Seymour et al. 2023; Burema et al. 2023), has prompted researchers to work on explainability techniques. Different taxonomies for such methods exist, e.g., (Guidotti et al. 2018), but one common classification is on the type of explanation generated (Molnar 2022). Model-based methods involve creating interpretable surrogate models, such as decision trees or linear models, which approximate the complex black box ML model (Ribeiro, Singh, and Guestrin 2018; Silva et al. 2020). Feature-based methods focus on pinpointing important features of the input, such as words in text or parts in an image, which contribute the most to the prediction (Ribeiro, Singh, and Guestrin 2016; Fong, Patrick, and Vedaldi 2019; Ancona 2017). Example-based methods provide explanations for a specific target outcome by deriving the importance of training samples (Ilyas et al. 2022; Garima et al. 2020; Koh and Liang 2017; Kwon and Zou 2022; Ghorbani and Zou 2019; Park et al. 2023), or provide a global overview of the model identifying representative examples (Yeh et al. 2018; Pruthi et al. 2020; Kim, Khanna, and Koyejo 2016; Gurumoorthy et al. 2019).
Example-based explainability offers several advantages. They are typically model-agnostic, and offer easy to understand explanations. More importantly, as they seek to discover a causal relationship between training examples and model behavior, they can assist in model debugging and data cleansing (Hara, Nitanda, and Maehara 2019), and be flexible for practitioners with different access (Richardson et al. 2023). However, they have two key limitations.
First, they don’t offer contrastivity (Nauta et al. 2023), which is key aspect in how humans understand decisions (Lipton 1990). While most methods can distinguish between supporters (aka proponents, helpful or excitatory examples), and opposers (aka opponents, harmful or inhibitory examples), they do not relate this information to ground truth labels (examples of class same as or different than predicted) or to the explanation intent (is the prediction correct/wrong, hard to tell). Contrastivity is the hallmark feature of counterfactual explanations (Wachter, Mittelstadt, and Russell 2017) and a major part of their appeal. Counterfactual explanations help users understand where the model’s decision boundary lies (Wachter, Mittelstadt, and Russell 2017), can offer algorithmic recourse (Karimi et al. 2022), and can audit a model for unfairness (Kavouras et al. 2023). However, by design, counterfactuals are imaginary instances, not necessarily plausible (Pawelczyk, Broelemann, and Kasneci 2020) or robust (Slack et al. 2021), and essentially offer feature-based explainability, revealing the important feature values contributing to the outcome.
@@
We propose a novel Antithetical, Intent-based, and Diverse Example-based (AIDE) explainability method, that offers contrastivity and is robust to outliers. At its core, AIDE is based on the concept of influence functions (Hampel 1974; Koh and Liang 2017). For a fixed target instance, the influence of a training sample is a score conveying its impact on the classifier’s outcome. Ideally, the influence is the change observed in the loss value for the target if the training sample was excluded from the training data. While influence scores can be estimated by methods, such as TraceIn (Garima et al. 2020) and Datamodels (Ilyas et al. 2022), we use the framework of the influence function approach (Koh and Liang 2017), termed IF, to efficiently compute influence scores.
To better understand AIDE’s contribution, we first showcase the issues that plague example-based explainability methods, taking IF as the representative—extensive qualitative and quantitative comparison with other methods is presented in the evaluation section. Consider a classifier that predicts whether short text messages are spam. Figure 1 shows that for the depicted target message, the prediction is spam. This is a correct prediction, and IF identifies the four most influential training samples at the top of Figure 1. We observe that explanations lack diversity, as they are highly similar to each other. More importantly, however, they lack contrastivity, as the user does not gain any insight about how the model decides what is spam and what not; all the user learns is that similar texts were labelled spam. The issue of susceptibility to outliers does not manifest in this example, mainly because the prediction is clearly correct. However, it manifests when, for example, it is not clear what the correct prediction should be, as in Figure 2.
@@
Figure 2: Explanations to clarify an ambiguous prediction.
AIDE is intent-aware. A critical review (Keane et al. 2021) identifies misalignments between the design goals of explainable AI (XAI) methods and the psychological and cognitive aspects of explainability. A key limitation they identify is that the majority of XAI methods neglect user intent during design and evaluation. Although mostly referring to counterfactual explanations, the authors highlight the importance of considering the user’s needs and goals while generating explanations. Moreover, the authors identify another common limitation in that explanations may not be plausible, which they define as not being relevant to the prediction being explained. AIDE explicitly addresses both limitations by offering intent-aware and relevant example-based explanations.
@@
AIDE outperforms state-of-the-art example-based methods. We perform an extensive quantitative and qualitative comparison against state-of-the-art methods for example-based explainability. Shapley-based approaches (Ghorbani and Zou 2019; Kwon and Zou 2022) were excluded as (a) they primarily aim to capture the overall contribution of training samples to the trained model (data valuation), and not geared for local explanations, and (b) they can be impractical for local explainability due to their high computational cost. The main conclusions drawn are as follows. Datamodels (Ilyas et al. 2022) approximate well the influence scores, but perform poorly for outlier targets. The principal reason is that Datamodels explain a class of models, and not a particular model. They thus fail to identify nuances picked up by a single model; while an unambiguous target will receive similar predictions by all models in the class, models will greatly differ in their predictions for an ambiguous target. TraceIn (Garima et al. 2020) is highly susceptible to outliers in the training data and performs poorly in tests of correctness and truthfulness. The reason lies in the way TraceIn estimates influence: it considers the difference in the total (training) loss when a training sample is included or not during checkpoints; outliers have high individual loss, contributing significantly to the total loss, and are thus awarded high importance. Regular IF are similarly affected by outliers in the training data. RelatIF (et al. 2020) seeks to address this problem, by penalizing samples that have high loss. However, these high-loss samples can at times be highly informative. For example, to explain a target instance that is ambiguous, it is often insightful to present those outlier training examples that are similar, so as to potentially uncover interesting labeling rules or protocols. In contrast, AIDE considers outliers as long as they are relevant to the target instance.
@@
Another method similar in principle to IF is TraceIn (Garima et al. 2020) that measures the influence of a training sample $X$ on a specific test sample $X_{0}$ as the cumulative loss change on $X_{0}$ due to updates from mini-batches containing $X$. They practically approximate this with TraceInCP, which considers checkpoints during training and sums the dot product of gradients at $X$ and $X_{0}$ at each checkpoint. Another unconventional work (Ilyas et al. 2022) fixes a test point to explain and samples a large number of subsets from the training set and trains models with each of these subsets. It then trains a linear model where the input will be $1_{S}i$ encoding of a subset and the output is the performance of the model trained on this subset for the test sample of interest. The weights of the linear model will represent the importance score of a training sample in the same position. To obtain a good result a huge number of intermediate models has to be trained on subsets, which is exhaustive, and thus a faster version of datamodels was proposed by (Park et al. 2023) and claimed to preserve almost the same accuracy. Since our focus is on the effectiveness of explanation we still use the original datamodels as a baseline.
@@
Evaluation of Explanations. While significant progress has been in developing explainability methods, there is a lack of standardized metrics for evaluating their effectiveness. The authors of (Doshi-Velez and Kim 2017) distinguish three types of evaluation strategies: application-grounded, human-grounded, and functionality-grounded. A profound study of functionality-grounded strategies by (Nauta et al. 2023), advocates twelve quantifiable properties that can be evaluated to assess the quality of explanations. They categorize the state-of-the-art metrics into twelve classes depending on which property the metric focuses on and what type of explanation is provided. The following properties are most relevant for local, example-based explanations: (1) Consistency and continuity both describe how deterministic the explanation is concerning identical and similar samples, assuming that these samples should have identical and similar explanations. In many works, this aspect is also referred to as the faithfulness (Jacovi and Goldberg 2020; Adebayo et al. 2018) of explanation and has gained popularity in the explainability domain. (2) Contrastivity is the ability of an explanation to interpret classes different than the prediction class. (3) Compactness is encoded in the size of an explanation as well as calculating a redundancy in the explanation. (4) Context describes how relevant the explanation is to the user needs. (5) Controlled synthetic Data check—Controlled Experiment: a synthetic dataset is developed with predetermined reasoning, ensuring that the predictive model aligns with this reasoning, as verified through metrics like accuracy. An assessment is done to check whether the explanation provided by the model corresponds to the same reasoning embedded in the data generation process, (Adebayo et al. 2020; Chen et al. 2018). In another work (Mothilal, Sharma, and Tan 2020), two general metrics for example-based explainability are proposed, diversity, and proximity.
User Study for Explainability. The field of user studies for the evaluation of explainable AI is multifaceted, encompassing diverse methodologies, types of explanations, and user groups. Previous research has delved into the impact of explanation formats, the role of system transparency, and the influence of individual differences on user understanding and trust. In the study conducted by (Rong et al. 2023), the investigation delves into the methodologies employed by human-computer interaction and AI researchers when conducting user studies in the realm of eXplainable AI (XAI) applications. A comprehensive literature review serves as the foundation for their exploration, wherein four prominent evaluation dimensions for explainability are highlighted.
  • The first dimension focuses on Trust, emphasizing the importance of establishing trust by ensuring the validity and sensibility of the model’s reasoning (Paleja et al. 2021).
  • Understanding, in the context of XAI, pertains to users’ comprehension of machine learning models. Objective understanding involves quantifiable aspects of comprehension, with studies revealing that measurable comprehension is enhanced by saliency maps, counterfactual explanations, and feature importance, albeit with disagreements on the impact of white versus black-box models. This understanding is influenced by factors such as data modality, proxy task choice, and interactivity (Hase and Bansal 2020). On the other hand, Subjective understanding involves users’ personal perception and interpretation of the AI system, with a prevailing trend of utilizing model explanations for improvement. Challenges arise with the “illusion of explanatory depth” (IOED), where users may exhibit overconfidence bias in their understanding of complex systems, leading to disagreements on the impact of explanations (Gentile, Donmez, and Jamieson 2023).
  • The concept of Usability in XAI refers to the effectiveness and user-friendliness of AI systems. It encompasses users’ perceptions of helpfulness, satisfaction, and other dimensions crucial for a positive and effective user experience (Nourani et al. 2022).
  • “Human-AI collaboration” denotes the interaction and cooperation between human users and AI systems to collectively enhance decision-making performance (Alufaisan et al. 2021). This collaboration is facilitated through users relying on explanations provided by AI systems to improve their accuracy in decision-making tasks.
@@
It can be shown (Cook and Weisberg 1982) that under certain conditions (second order differentiability and convexity of the loss function) the derivative of ${\bm{\theta}}^{*}$ can be expressed:
@@
where ${\bm{H}}_{{\bm{\theta}}^{*}}$ is the Hessian matrix of the objective $L({\mathcal{S}},{\bm{\theta}}^{*})$ calculated at ${\bm{\theta}}^{*}={\bm{\theta}}^{*}_{0}$.
@@
To explain the prediction for ${\bm{z}}_{t}$, we use Equation 5 to compute the influence of each training example ${\bm{z}}$, which can be done efficiently as suggested in (Koh and Liang 2017). The IF explanation for the prediction for ${\bm{z}}_{t}$ consists of the top-$k$ training examples with the highest influence.
@@
Existing approaches for influence-based explainability (Koh and Liang 2017; et al. 2020) compile an explanation as a set of highly influential training examples. We claim that other aspects, besides high influence, are also important. Specifically, AIDE creates explanations that contain training examples with negative influence, considers their labels, their proximity to the test instance, and their diversity.
@@
Proof.
Note that the model assigns the same probability, $p=p_{\bm{\theta}}^{*}({\bm{x}}_{t})$, to ${\bm{z}}_{t}$ and ${\bm{z}}_{t}^{\prime}$ belonging in the positive class. Let $h({\bm{z}}_{t})=\left.\frac{\mathrm{d}\ell({\bm{z}}_{t},{\bm{\theta}}^{*})}{\mathrm{d}p_{\bm{\theta}}^{*}({\bm{x}}_{t})}\right|_{p_{\bm{\theta}}^{*}({\bm{x}}_{t})=p}$ be the derivative of the loss function with respect to the probability of ${\bm{z}}_{t}$ being positive. Then using the chain rule we can rewrite the gradient of the loss of ${\bm{x}}_{t}$ as:
${\bm{g}}({\bm{z}}_{t})=\left.\nabla_{{\bm{\theta}}^{*}}\ell({\bm{z}}_{t},{\bm{\theta}}^{*})\right|_{{\bm{\theta}}^{*}={\bm{\theta}}^{*}_{0}}=h({\bm{z}}_{t})\left.\nabla_{{\bm{\theta}}^{*}}p_{\bm{\theta}}^{*}({\bm{x}}_{t})\right|_{{\bm{\theta}}^{*}={\bm{\theta}}^{*}_{0}}.$
Substituting ${\bm{g}}({\bm{z}}_{t})$ into the influence of any example ${\bm{z}}$, we get:
$I({\bm{z}},{\bm{z}}_{t})=h({\bm{z}}_{t})\left.\nabla_{{\bm{\theta}}^{*}}^{\intercal}p_{\bm{\theta}}^{*}({\bm{x}}_{t})\right|_{{\bm{\theta}}^{*}={\bm{\theta}}^{*}_{0}}{\bm{H}}^{-1}_{{\bm{\theta}}^{*}}{\bm{g}}({\bm{z}})=h({\bm{z}}_{t})\hat{I}({\bm{z}}),$
where we have named $\hat{I}({\bm{z}})$ the part of influence that does not depend on ${\bm{z}}_{t}$. So we get:
$I({\bm{z}},{\bm{z}}_{t})=\frac{h({\bm{z}}_{t})}{h({\bm{z}}_{t}^{\prime})}I({\bm{z}},{\bm{z}}_{t}^{\prime}).$
The derivative of the logistic loss for ${\bm{z}}_{t}$ w.r.t. the prediction probability is $h({\bm{z}}_{t})=\frac{p-y_{t}}{p(1-p)}$. Similarly, for ${\bm{z}}_{t}^{\prime}=({\bm{x}}_{t},1-y_{t})$ we get $h({\bm{z}}_{t}^{\prime})=\frac{p-1+y_{t}}{p(1-p)}$. Thus, the ratio $h({\bm{z}}_{t})/h({\bm{z}}_{t}^{\prime})$ becomes $-\frac{1-p}{p}$ when $y_{t}=1$, and takes the inverse value when $y_{t}=0$. ∎
@@
Proximity. Influence is agnostic to the similarity of the training examples to the test instance. As noted (et al. 2020), there may exist outliers and mislabeled training examples that can exhibit high magnitude influence scores. Such examples are often globally influential, i.e., they are influential for many test instances, just because they are unusual. These are rarely useful as an explanation, and (et al. 2020) proposes to normalize the influence of an example with their global influence. Nonetheless, in certain cases these outliers are extremely useful, e.g., when explaining another outlier.
@@
Diversity. Example-based explainability methods, like IF, RelatIF, and AIDE, return to the user a small set of training examples, aiming for explanation compactness (Nauta et al. 2023). It is thus important that the set of examples avoids redundancy. AIDE, in contrast to prior work (Koh and Liang 2017; et al. 2020), considers the diversity of the explanation set. Assuming an internal representation of training examples and an appropriate similarity measure $\operatorname{sim}$, we define diversity of a set ${\mathcal{E}}$ of training examples as $D({\mathcal{E}})=1-\frac{1}{|{\mathcal{E}}|(|{\mathcal{E}}|-1)}\sum_{{\bm{z}},{\bm{z}}^{\prime}\in{\mathcal{E}}}\operatorname{sim}(\hat{{\bm{x}}},\hat{{\bm{x}}^{\prime}})$.
@@
Support. It comprises examples with positive influence and with the same label as the test instance. They play a positive role in the prediction and resemble the test instance: “You get the same outcome with these”.
@@
To select the appropriate examples for each quadrant, we perform a series of steps. After partitioning the training examples in the four quadrants, we select only examples with high magnitude. We use the Interquartile Range (IQR) method, (Agresti and Franklin 2005), to keep examples with positive influence above $Q_{3}+\lambda IQR$, and to keep examples with negative influence below $Q_{1}-\lambda IQR$, where $Q_{1}$ and $Q_{3}$ are the first and the third quartiles of the influence distribution, $IQR=Q_{3}-Q_{1}$, and $\lambda$ is a coefficient that controls the number of high-magnitude influential points, and is empirically determined. After this filtering, we end up with a candidate set ${\mathcal{S}}_{q}$ of training examples for each quadrant.
@@
In our experiments, we used two datasets: the SMS Spam dataset [^2] https://www.kaggle.com/datasets/uciml/sms-spam-collection-dataset, which comprises a collection of text messages labeled as either spam or non-spam (ham), commonly used for text classification and a derivative dataset with pictures of dogs and fish extracted from Imagenet [^3] https://www.image-net.org/. For the spam classification task, we employed the BERT-base pre-trained word embedding model and incorporated two sequential layers to capture the specific characteristics of our data. Regarding the image classification task, we utilized a pre-trained InceptionV3 model removing the output layer and appending sequential layers to learn the peculiarity of our task. All the baselines were implemented with instructions given in their papers and GitHub repositories. While running the greedy algorithm for the weighted sum of AIDE properties, we determined the optimal weights denoted as $\alpha$, $\beta$, and $\gamma$, through iterative refinement and empirical exploration, for each specific task. Fine-tuning these weights allowed us to achieve an optimized sampling strategy and enhance the effectiveness of the explanations provided by AIDE. In the case of text classification, we assigned a higher weight to proximity, as the influence was already considered through the use of IQR and the selection of the most influential samples. Furthermore, the similarity between textual data, for which BERT embeddings were generated, was found to be well-captured. Conversely, in image classification, the similarity metric did not always accurately reflect proximity. Therefore, we emphasized the influence factor by assigning a higher weight to it. To prevent the selection of identical or highly similar samples multiple times, we introduced a similarity threshold. This ensures that such samples are not chosen repeatedly, thus a smaller coefficient was assigned to promote diversity. The coefficient for IQR was set to $\lambda=3$ in all cases. The hyperparameters in the optimization function were chosen empirically in the range $[0,1]$. We observed that the diversity weight $\gamma$ does not affect the quality of the explanations that much, as long as it was nonzero; we set it to $\gamma=0.5$ in all experiments. The other two hyperparameters control the presence of outliers in the explanations; higher values of $\beta$ suppress outliers by giving more weight to training examples that are similar and have high enough influence. We settled to $\alpha=0.2$ and $\beta=0.8$ for all experiments. The baseline methods that we will compare AIDE to are IF, RelatIF , Datamodels , and TraceIn.
The documented code for replicating our experiments and the associated data can be accessed in the GitHub repository [^4] https://anonymous.4open.science/r/aide-3E06/README.md. Our implementation utilizes the implementation of influence functions from the repository [^5] https://github.com/alstonlo/torch-influence.git under a licence [^6] https://github.com/alstonlo/torch-influence/blob/main/LICENSE.txt. The datasets employed in our experiments include the Spam collection dataset [^7] https://creativecommons.org/licenses/by/4.0/legalcode, and a derivative dataset of ImageNet  [^8] https://www.image-net.org/update-mar-11-2021.php.
We ran our experiments in a server with the following specifications: Dell R6415, 256GB RAM, 16 cores (AMD 7281), 240GB SSD, 8TB HDD, 2x25 Gbit.
Complexity Analysis. To assess the time complexity, we first consider the time required to compute the influence scores for a given prediction. Using the techniques suggested in (Koh and Liang 2017), this takes $O(np)$ time, where $n$ represents the size of the training set, and $p$ is the number of learned parameters. Subsequently, AIDE employs a greedy-based iterative algorithm for selecting the $k$ best training samples in each quadrant, which takes $O(nk)$ time. The cost to eliminate training samples based on the IQR of their influence is linear in $n$ and is excluded. Overall the running time of AIDE is $O(np+nk)$, and is dominated by the time required to compute influence scores.
@@
Figure 3: Continuity in terms of explanation similarity vs. instance pair similarity in spam dataset.
(a) IF
(b) RelatIF
(c) AIDE
(d) DM
(e) TracIn
Figure 4: Continuity in terms of explanation similarity vs. instance pair similarity in image dataset.